OpenStudy (anonymous):

solve for x log7 10-1/3log7 x+log7 2

OpenStudy (rajee_sam):

\[x \log_{7} 10 -\frac{ 1 }{ 3 }\log_{7} x + \log_{7} 2\] Is this your problem?

OpenStudy (rajee_sam):

my bad the x is not there in the front?

OpenStudy (rajee_sam):

also if you want to solve for x it should be = to something. What is missing there?

OpenStudy (rajee_sam):

\[\log_{7} 10 - \frac{ 1 }{ 3 }\log_{7} x + \log_{7} 2 = ??\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry its equal to the log of 7 2

OpenStudy (rajee_sam):

OK so your problem is\[\log_{7} 10 - \frac{ 1 }{ 3 }\log_{7} x = \log_{7} 2\]

OpenStudy (rajee_sam):

do you know the logarithmic rules?

OpenStudy (anonymous):

yes thank you so much can u help me and explain?

OpenStudy (anonymous):

no my teacher doesnt teach that much

OpenStudy (anonymous):

i have to use the book and im struggling

OpenStudy (rajee_sam):

\[\log_{a} A + \log_{a} B = \log_{a} (A . B)\]

OpenStudy (rajee_sam):

\[\log_{a} A - \log_{a} B = \log_{a} \frac{ A }{ B }\]

OpenStudy (anonymous):

umm can you break it down into words too?

OpenStudy (rajee_sam):

Log of A to the base a + Log of B to the base a = Log (A times B) to the base a Note: The base should be the same for you to do that Similarly for Log A to the base a - Log B to the base a = Log ( A Divided by B) to the base a. eg Log 2 + Log 3 = Log ( 2x 3) = Log 6, they both should have the same base. For convenience I did not write the base.

OpenStudy (rajee_sam):

Log 6 - Log 3 = Log (6/3) = Log 2

OpenStudy (anonymous):

ok so i would use the first equation

OpenStudy (rajee_sam):

there is 3rd rule

OpenStudy (rajee_sam):

\[m \log_{a} A = \log_{a} A ^{m}\]

OpenStudy (rajee_sam):

eg. 3 Log 2 = log 2³ = log 8

OpenStudy (rajee_sam):

pl. write down these rules somewhere all in one place so you can refer to it whenever you need them

OpenStudy (anonymous):

ok thank you soo soo mcuh one last thing could you solve an equation like mine so i have an example

OpenStudy (rajee_sam):

4th rule\[\log_{a} a = 1 ; ( example.) \space \log_{5} 5 = 1\]

OpenStudy (rajee_sam):

fifth rule

OpenStudy (rajee_sam):

Log 1 to any base = 0

OpenStudy (rajee_sam):

All these five rules write them down Now we will use these rules to solve your problem

OpenStudy (anonymous):

ok i wrote them

OpenStudy (rajee_sam):

now in your problem ( in any problem ) first thing you check is , Is all the bases the same. Here in your problem all logs on either side have the same base which is 7. So from now on for this problem if I do not write the base it is understood that the base is 7 ok.

OpenStudy (anonymous):

ok

OpenStudy (rajee_sam):

Now we have 3 logs in the problem . The first one Log 10 we cannot simplify any further. So we keep it The second one 1/3 log x. This is of the form m log A, m = 1/3 and A = x. Now tell me which rule do you use to rewrite it?

OpenStudy (anonymous):

3rd rule

OpenStudy (rajee_sam):

yes . can you rewrite it?

OpenStudy (anonymous):

ill try just a sec

OpenStudy (anonymous):

\[-1/3\log _{7}x=\log _{7}x ^{-1/3}\]

OpenStudy (anonymous):

?

OpenStudy (rajee_sam):

in a way yes. but I will leave the negative sign alone and just keep a positive 1/3 for the power.

OpenStudy (rajee_sam):

\[\log_{7} 10 - \frac{ 1 }{ 3 }\log_{7} x = \log_{7} 10 - \log_{7} x ^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

ok so subtraction becomes divison?

OpenStudy (rajee_sam):

absolutely

OpenStudy (rajee_sam):

remember we are just on our left hand side of the equation, we have not gone to the right yet

OpenStudy (rajee_sam):

so let us go ahead and further simplify these 2 logs into one log using rule no. 2

OpenStudy (anonymous):

ooh ok

OpenStudy (rajee_sam):

can you do it?

OpenStudy (anonymous):

im gonna try

OpenStudy (rajee_sam):

good I 'll wait

OpenStudy (anonymous):

\[\log _{7}10-\log _{7}x=\log _{7}\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

o.o

OpenStudy (rajee_sam):

you can just type it out for now

OpenStudy (anonymous):

ok

OpenStudy (rajee_sam):

you try the tools for this later

OpenStudy (anonymous):

k

OpenStudy (anonymous):

cool thanks for the medal does that mean i got that part right sorta?

OpenStudy (rajee_sam):

no... I just like the way you try things out. but that was way wrong

OpenStudy (anonymous):

oooh can u show me how then?

OpenStudy (rajee_sam):

\[\log_{7} 10 - \log_{7} x ^{\frac{ 1 }{ 3 }} = \log_{7} \frac{ 10 }{ x ^{\frac{ 1 }{3 }} }\]

OpenStudy (rajee_sam):

Log A - Log B = Log (A/B), A = 10 and B = x^1/3

OpenStudy (rajee_sam):

now that is your Left hand side of the equation. remember your right hand side was log 2 to the base 7?

OpenStudy (anonymous):

ya

OpenStudy (rajee_sam):

Ok now since both sides have one log and both are same base. We can drop the log like this. if Log A = Log B then A = B

OpenStudy (anonymous):

what?how are they equal

OpenStudy (rajee_sam):

That is the rule no. 6 I\[\log_{a} A = \log_{a} B ; then \space A = B\]

OpenStudy (anonymous):

ok i understand that but how are they equal

OpenStudy (rajee_sam):

we still have to solve for x

OpenStudy (anonymous):

ok

OpenStudy (rajee_sam):

\[\log_{7} (\frac{ 10 }{ x ^{\frac{ 1 }{ 3 }} }) = \log_{7} 2\]means\[\frac{ 10 }{ x ^{\frac{ 1 }{ 3 }} } = 2\]

OpenStudy (rajee_sam):

Now cross multiply

OpenStudy (rajee_sam):

10/2 = x^1/3

OpenStudy (rajee_sam):

5 = x^1/3

OpenStudy (rajee_sam):

\[x^{\frac{ 1 }{ 3 }} = 5\]

OpenStudy (anonymous):

125?

OpenStudy (rajee_sam):

now cube both sides

OpenStudy (rajee_sam):

yes

OpenStudy (anonymous):

yay now i feel sorta smart

OpenStudy (rajee_sam):

x = 125

OpenStudy (rajee_sam):

So use the rules I gave you and you will always use a couple of them and solve your x

OpenStudy (rajee_sam):

well done

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\left[\begin{matrix}x+y= & 9\\ 4x-5y= & 0\end{matrix}\right]\]

OpenStudy (anonymous):

hw do u inverse this matrixs

OpenStudy (rajee_sam):

could you pl. post it in a new window?